### v1

# Fine-Grained Vision-Language Modeling for Multimodal Training Assistants in Augmented Reality

Abstract
[…]
VLM are essential for enabling AI-powered smart assistants to interpret and reason in multimodal environments. However, their application in augmented reality (AR) training remains largely unexplored. In this work, we introduce a comprehensive dataset tailored for AR training, featuring systematized vision-language tasks, and evaluate nine state-of-the-art VLMs on it. Our results reveal that even advanced models, including GPT-4o, struggle with fine-grained assembly tasks, achieving a maximum F1 score of just 40.54% on state detection. These findings highlight the demand for enhanced datasets, benchmarks, and further research to improve fine-grained vision-language alignment. Beyond technical contributions, our work has broader social implications, particularly in empowering blind and visually impaired users with equitable access to AI-driven learning opportunities. We provide all related resources, including the dataset, source code, and evaluation results, to support the research community.
1 Introduction
Training assistants are in high demand in industrial assembly tasks, including furniture production (You et al., 2022), car manufacturing (Bellalouna et al., 2020), and industrial product assembly (Funk et al., 2017). As these tasks often require high levels of precision and cognitive effort, augmented reality (AR) has emerged as a promising technology to support workers, providing a multimodal, interactive, user-centric environment (Hou & Wang, 2013; Botto et al., 2020; Dalim et al., 2020). AR training systems have been extensively studied for their ability to enhance user experience and task efficiency (Büttner et al., 2020; Blattgerste et al., 2021). However, most existing research focuses on interaction design and usability studies, rather than the development of intelligent, instruction-following assistants that can understand and process complex manual-based assembly tasks in AR environments.
[…]
VLM offer a powerful solution for bridging visual and textual modalities by effectively interpreting and integrating visual inputs with corresponding textual instructions. Traditional computer vision studies have explored various purely visual tasks, such as action segmentation, recognition, anticipation, and object and state detection (Damen et al., 2018; Miech et al., 2020; Wang et al., 2023b; Gao et al., 2024). Some studies focus on analyzing procedural activities in videos without fixed instructions (Tang et al., 2019; Sener et al., 2022), while others incorporate textual instructions to enhance comprehension. These approaches include using instructional queries for response generation(Li et al., 2023a; Padmakumar et al., 2021), narration-video alignment(Zhukov et al., 2019), and text-to-video retrieval(Miech et al., 2019). However, none of these studies explicitly address instruction-following in manual assembly tasks, where models must accurately interpret step-by-step procedural instructions and track fine-grained object states within a structured pipeline. The recent advancements in large language models and vision-language models have made multimodal reasoning more accessible, with models such as GPT-4(Achiam et al., 2023), LLaVa (Liu et al., 2023), QWEN-VL (Bai et al., 2023), BLIP2 (Li et al., 2023c), and MiniGPT-v2 (Chen et al., 2023). The core idea is to excel in multimodal tasks by integrating visual encoders with fundamental LLMs. Some early-stage explorations have attempted to apply VLMs in AR to align visual contexts with language representations to showcase the capabilities of LLM agents (Pei et al., 2024). However, these efforts have primarily focused on coarse-grained visual features, such as detecting objects and their coordinates, rather than understanding complex procedural assembly tasks in detail. A key challenge in AR-based assembly training is the fine-grained interpretation of object states and spatial relationships. Figure 1 illustrates this limitation: when prompted with an image and text instructions for LEGO assembly, GPT-4o fails to correctly identify an incorrect assembly state, where a newly added LEGO part (in green) is misplaced. This example highlights a crucial weakness – current VLMs struggle with precise state detection and fine-grained assembly verification, making them unreliable for industrial AR training scenarios.
Figure 1: An example of incorrect state detection by GPT-4o during LEGO assembly. Given an image and text instructions, the model fails to accurately recognize an incorrectly placed LEGO part (highlighted in green), demonstrating the challenge of fine-grained vision-language alignment in assembly tasks.
To address these challenges, we introduce a comprehensive vision-language benchmark and dataset designed for AR-based assembly training. Specifically, we chose LEGO brick assembly in AR as a simplified simulation and created a synthetic multimodal dataset to address vision-language challenges, i.e., complex scene understanding, object detection, and state detection. We developed a unified, fine-grained VLM architecture capable of handling a broad range of vision-language tasks in AR and validated the feasibility of nine prevailing models as benchmarks. Notably, this research is designed for various manual-guided assembly assistance in a multimodal environment, though we specifically demonstrated it using LEGO assembly. We list contributions below: (1) Developing a unified VLMs-based architecture to benchmark fine-grained instruction-following capabilities in AR training. (2) Evaluating nine prevailing VLMs, both without and with fine-tuning, on a synthetic multimodal dataset tailored for AR-based LEGO assembly. (3) Creating and releasing a dataset containing systematic vision-language tasks for AR training, along with a method for synthesizing high-quality multimodal data.
2 Fine-grained Vision-Language Modeling
2.1 Task definition
As illustrated in Figure 2, we investigate manual-guided LEGO assembly in AR and define the following three vision-language tasks: (T1) Scene Understanding. Given the current step’s image $V_{t}$ and the task-specific query $Q_{t}$, the model outputs a response containing the scene description for the current assembly step $I_{t}$. $Q_{t}$ is the task description contextualized with the previous step’s textual instruction $I_{t-1}$. It assesses the model’s ability to accurately describe the current assembly step from an instruction manual, including but not limited to generating object representations, their properties, and the assembly procedure. (T2) Object Detection. Given a seeking step’s image $V_{t^{\prime}}$ ($t^{\prime}=t-1$) and task-specific query $Q_{t^{\prime}}$, the model outputs a response containing the positional text $P_{t^{\prime}}$, which includes identified object and its coordinate formatted as “<Object> <Xleft> <Ytop> <Xright> <Ybottom>”. The query $Q_{t^{\prime}}$ is the task description contextualized with the corresponding textual instruction $I_{t^{\prime}}$. It assesses the model’s ability to accurately identify and locate objects in the scene, ensuring they are positioned correctly for the next assembly action. (T3) State Detection. Given an assembly step’s image, either the original manual image $V_{t}$ or a generated negative sample $\tilde{V}_{t}$, and task-specific query $Q_{t}$, the model outputs a response indicating a correct or incorrect state. It evaluates the model’s ability to verify the accuracy of the assembly progression, determining if the assembly action has been correctly completed. These tasks have been crafted to rigorously test and demonstrate the capabilities of VLMs in an AR training, focusing on their ability to understand and interpret complex multimodal inputs.
Figure 2: An illustration of the proposed three VLM tasks, highlighting the core capabilities required by the one-for-all VLM architecture for LEGO assembly based on the procedural instruction manual.
2.2 One-for-all architecture
We investigated existing VLMs and derived a universal architecture for the proposed tasks, as illustrated in Figure 3. Given an image, either from manuals or real-world scenes, and a task-specific query, a VLM based on this architecture generates a textual response as the output. This architecture integrates a vision encoder, a large language model, and a vision-language projector as core modules, along with a task-specific query for task adaptation. Specifically, the vision encoder processes visual inputs, the large language model interprets and generates textual information, and the vision-language projector aligns visual and textual data for seamless task execution. The query can be customized using the following key components: (1) Task-specific token:Special tokens such as “[grounding]”, “[object]”, and “[state]” are introduced for task T1, T2, T3, respectively, to enhance task focus and accuracy. (2) Instruction:Relevant manual instructions or task directives are incorporated to provide the model with context, ensuring responses align with task requirements. (3) Format-directive:A directive to clarify the expected output format, ensuring outputs are precise and directly applicable.
Figure 3: The architecture for one-for-all vision-language modeling. It integrates a vision encoder, a large language model, and a vision-language projector as core functional modules, along with a task-specific query for task adaptation.
3 Dataset Creation
3.1 Manual Crawling and Scene Data Matching
This subsection outlines the procedure for creating the dataset for task T1. First, we collected 65 instruction manuals from the official LEGO website [^1] https://legoaudioinstructions.com/instructions, designed to help blind and visually impaired users assemble LEGO sets accurately. Each instruction manual comprises several elements: (1) step-by-step textual instructions; (2) corresponding image for each step; and (3) tags provided by LEGO include instruction and image tags. Then, we split a full manual into several assembly sessions, each session includes two types of steps: object seeking and object assembly. Each seeking step is followed by an assembly step , forming a pair. Next, we iterated through all sessions to generate the scene understanding dataset, $D_{T1}=\{(Q_{t},V_{t},A_{t})\}|_{t=0}^{|D_{T1}|}$, where each element is a triplet of (query, image, text), and the query $Q_{t}$ is constructed by filling the query template with the task token “[grounding]” and the previous instruction $I_{t-1}$ as context.
3.2 Object Position Inference
This subsection outlines the procedure for creating the dataset for task T2, as illustrated in Figure 6.
First, we iterated through all object-seeking steps, where users are asked to find specific objects. Second, we iterated all images to generate the object detection dataset, $D_{T2}=\{(Q_{t^{\prime}},\tilde{V}_{t^{\prime}},A_{t^{\prime}})\}|_{t^{\prime}=0}^{|D_{T2}|}$, where each element is a triplet of (query, image, text) for a seeking step $t^{\prime}=t-1$. The query $Q_{t^{\prime}}$ is constructed by filling the query template with the task token “[object]” and the current instruction $I_{t^{\prime}}$. The response $A_{t^{\prime}}$ contains a positional text $\tilde{P}_{t^{\prime}}$ formatted as “<Object> <Xleft> <Ytop> <Xright> <Ybottom>”. This is initially generated by querying the image using MiniGPT-v2 (Chen et al., 2023), the state-of-the-art model at the time of this work. The composite image $\tilde{V}_{t^{\prime}}$ is created by combining the current image ${V}_{t^{\prime}}$ with three randomly sampled images. Last, the initial coordinates are adjusted to fit the composite image.
3.3 Variant State Generation
This subsection outlines the procedure for creating the dataset for task T3, as illustrated in Figure 7. First, we iterated through all object-assembly steps, along with their corresponding previous steps, where users are trained to assemble the parts identified in the prior steps. Second, we conducted part segmentation by detecting the boundary color and segmenting the part to be assembled. In each assembly session, the objects to be assembled in the current step are tagged in the previous step, with their boundaries highlighted in distinct colors that vary between sessions. The highlighted colors are selected based on predominant color detection using K-means clustering [^2] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html, followed by Hue filtering (Chu et al., 2014) and manual correction for accuracy. Third, we added bounding boxes to the segmented objects. Fourth, we applied a natural perturbation to the part to be assembled by randomly shifting it within the background box, ensuring it moves by at least 5%. We treated each state from the manual as a positive sample and generated three variant states as negative samples. Then, we constructed the state detection dataset, $D_{T3}=\{(Q_{t},V_{t},A_{t},\tilde{V}_{t},\tilde{A}_{t},\tilde{V}_{t}^{{}^{\prime}},\tilde{A}_{t}^{{}^{\prime}},\tilde{V}_{t}^{{}^{\prime\prime}},\tilde{A}_{t}^{{}^{\prime\prime}})\}|_{t=0}^{|D_{T3}|}$, where the query $Q_{t}$ is constructed by filling the query template with the task token “[state]”, the current instruction $I_{t}$, and the previous instruction $I_{t-1}$ as context. The variant states $\{\tilde{V}_{t},\tilde{V}_{t}^{{}^{\prime}},\tilde{V}_{t}^{{}^{\prime\prime}}\}$ are natural pertubations of $V_{t}$. The responses $A_{t}$ and $\{\tilde{A}_{t},\tilde{A}_{t}^{{}^{\prime}},\tilde{A}_{t}^{{}^{\prime\prime}}\}$ indicate the correct and incorrect statuses of the assembly state, respectively.
3.4 Dataset Statistics
We summarize the statistics of LEGO-VLM dataset in Table 1. It is generated from 65 LEGO instruction manuals and divided into 397 sessions, covering 5,612 scenes, 4,784 objects, and 2,716 states. It consists of 5,614 instruction steps, with each step containing an image and corresponding textual instructions. Out of these, 4,814 steps focus on object seeking, while 5,614 steps involve parts assembly, covering 3,172 states with 222 distinct boundary colors. Overall, the dataset includes 35,612 vision-language data samples: 5,612 for T1, 19,136 for T2, and 10,864 for T3, respectively.
Manual Session Scene Object State Step Sample
Overall Identification Assembly Overall T1 T2 T3
65 397 5,612 4,784 2,716 10,428 4,814 5,614 35,612 5,612 19,136 10,864
Table 1: Statistics of LEGO-VLM dataset.
4 Experimental Setup
4.1 Benchmarks
We benchmark the nine prevailing VLMs: (1) mPLUG-OWL2 (Ye et al., 2024)replaces attention with a modality adapter in a LLM decoder. (2) BLIP2 (Li et al., 2023c)uses a two-stage pre-trained Q-Former between an image encoder and a LLM to bridge the vision-language modality gap. (3) LLaVa (Liu et al., 2023)combines a visual CLIP encoder and a language decoder Vicuna for general-purpose visual language understanding. (4) Qwen-VL (Bai et al., 2023)connects a LLM with a visual encoder using position-aware vision-language adapter towards fine-grained visual understanding. (5) InstructBLIP (Dai et al., 2023)explores general-purpose vision-language instruction tuning based on the pretrained BLIP2. (6) MiniGPT-v2 (Chen et al., 2023)links a frozen ViT visual encoder with Llama-2 via a projection layer, applicable for diverse tasks via task-specific multimodal instructions. (7) Otter (Li et al., 2023a)is tuned on the OpenFlamingo, conditioning the language model on images for multi-modal perception and reasoning. (8) MiniGPT-4 (Zhu et al., 2023)integrates a frozen visual ViT&Q-Former encoder and Vicuna via a projection layer, unlocking advanced multimodal capabilities like GPT-4. (9) GPT-4o [^3] https://openai.com/index/hello-gpt-4o/is a widely used commercial model accessible via APIs.
4.2 Evaluation Metrics
[…]
5 Outcomes
5.1 Benchmark Results
We compare the performance of nine prevailing VLMs on the proposed LEGO-VLM dataset, without and with finetuing, as shown in Table 2.
Model T1: Scene understanding T2: Object detection T3: State detection
F1-Theme $\uparrow$ ROUGE $\uparrow$ BLEU $\uparrow$ F1-Object $\uparrow$ IOU $\uparrow$ F1-State $\uparrow$ FPR $\downarrow$
PEFT (LoRA) w/o w/ w/o w/ w/o w/ w/o w/ w/o w/ w/o w/ w/o w/
mPLUG-OWL2 Ye et al. (2024) 23.16 29.70 25.23 32.75 3.04 8.46 77.94 93.05 14.23 34.88 36.82 15.00 63.41 35.00
BLIP2 Li et al. (2023c) 27.85 32.65 29.55 40.35 6.40 12.50 77.35 84.48 34.25 40.57 24.71 28.50 100.00 94.87
LlaVA Liu et al. (2023) 32.10 35.88 13.63 42.97 0.99 17.73 54.39 72.32 $\emptyset$ 60.98- 25.04 30.00 99.41 50.00
QWEN-VL Bai et al. (2023) 34.84 37.00 29.57 39.08 5.19 13.13 78.56 89.15 25.60 30.08 39.77 39.53 97.99 96.82
InstructBLIP Dai et al. (2023) 32.85 32.85- 29.20 36.87 4.91 11.52 79.76 98.16 $\emptyset$ 47.20- 0.00 0.02 1.22 0.00
MiniGPT-v2 Chen et al. (2023) 33.06 37.52 34.72 32.56 9.81 8.28 84.95 85.91 26.98 25.94 36.76 38.64 60.42 80.72
Otter Li et al. (2023a) 11.49 12.12 1.28 72.39 $\emptyset$ 35.33 75.88
MiniGPT-4 Zhu et al. (2023) 34.11 15.05 1.88 87.09 30.20 37.19 67.37
GPT-4o 44footnotemark: 4 25.81 18.67 2.00 66.67 21.68 40.54 43.06
Table 2: Benchmarking vision-language models on LEGO-VLM dataset, without (/wo) and with (/w) PEFT using LoRA. The bold font indicates the highest score in each column. Symbols $\uparrow$ and $\downarrow$ denote that higher and lower values are better, respectively. Symbol “–” indicates the model is not applicable for fine-tuning. Symbol $\emptyset$” denotes a meaningless zero as the model fails to generate output as instructed. The superscripts “”, “”, and “-” indicate an increase, decrease, or inapplicability in the evaluation score after fine-tuning, respectively.
First of all, existing models struggle with fine-grained assembly tasks in AR. While the F1-Object score for task T2 using the fine-tuned InstructBLIP is high at 98.16%, the intersection over union (IOU) is only 47.20%, indicating minimal overlap between the predicted object positions and the ground truth, which is still unsatisfactory. Even the commercial model GPT-4o achieves only 66.67% for the F1-Object score and 21.68% on task T2. This might also caused by the difficulty of understanding scenes. For example, fine-tuned LlaVA shows the best overlap with reference instructions, scoring 42.97% on ROUGE and 17.73% on BLEU. However, the fine-tuned MiniGPT-v2, despite being the top performer in theme entity identification, achieves only 37.52%. This indicates that understanding and generating theme entities, such as LEGO parts and their properties, remains a significant challenge. For state detection in task T3, the fine-tuned InstructBLIP achieves a 0.00% of FPR, meaning it almost perfectly avoided incorrectly classifying negative instances as positive. However, the F1-State score is just 0.02%, indicating a significant failure in identifying the correct states. These results highlight the challenges of the proposed fine-grained vision tasks, reinforcing the need to advance this area and its resources as a valuable research topic.
Second, open-resource VLMs with fine-tuning generally outperform or are comparable to commercial models. For one hand, top-performing open-resource VLMs are more informative with identified accurate entities. For example, MiniGPT-v2 with fine-tuning achieves an 11.71% higher F1-Theme score on task T1; InstructBLIP with fine-tuning achieves a 31.48% higher F1-Object score, while QWEN-VL shows only a 1% decrease, compared with the commercial GPT-4o. This may be due to the proposed dataset enhancing the model’s understanding of domain-specific knowledge, such as LEGO parts and their properties. On the other hand, top-performing open-resource VLMs can generate assembly instructions that closely align with the provided manuals. For example, the fine-tuned LlaVA achieves ROUGE and BLEU scores that are 2.30 and 8.87 times higher than those of GPT-4o, respectively. Besides, fine-tuning generally improves the performance of VLMs, as indicated by the results marked with the superscript “”. This highlights the significant difference between the proposed dataset and those used to train general commercial models.
Last, task difficulty varies significantly, with the biggest challenge being the alignment of positional information in images with the corresponding textual information from the query. Specifically, in comparison to task T1, tasks T2 and T3 present greater challenges in instruction following. For example, several VLMs (i.e., LlaVA, InstructBLIP, Otter), without fine-tuning, fail to follow instructions to generate positional information for evaluating IOU, indicated by the symbol $\emptyset$.” Additionally, all evaluated VLMs have F1-State scores below 50%, indicating that their predictions are even worse than random guessing. One potential reason task T3 is complex, and relies on task T2 is that it requires positional information to detect states, as well as insights from task T1 to understand the LEGO components and their relationships in the scene. We will explore this in future work, as this study primarily focuses on proposing the tasks rather than complex modeling.
5.2 Data Quality Assessment
To ensure the quality of the generated data for tasks T2 and T3, we sampled 100 data points and conducted quality assessments for each task. We added bounding boxes based on the coordinates and asked three annotators to evaluate the generated data quality, focusing on coordinates, entities, and negative samples, based on the following criteria: (1) Entity disambiguitymeasures how clearly a target entity is relevant to the scene in T2, with scores of 0, 1, and 2 indicating low, medium, and high disambiguation, respectively. (2) Boundary precisionmeasures how accurately a bounding box encloses a target object in T2, with scores of 0, 1, and 2 indicating low, medium, and high precision, respectively. (3) State relevancemeasures whether the generated parts in an image are a relevant variation of the original image in T3. (4) State identifiabilitymeasures whether the generated parts in an image are a recognizable variation of the original image in T3.
Figure 4: Results of the quality assessment.
As shown in Figure 4, the average entity disambiguation and boundary precision scores are 1.74 and 1.25 out of 2 with the Kappa values [^5] https://www.statsmodels.org/dev/generated/statsmodels.stats.inter_rater.fleiss_kappa.html of 0.60 and 0.61, respectively. This suggests that annotators reached a moderate agreement on the clarity of the evaluated objects in relation to the corresponding scenes. The average state relevance and identifiability scores are 0.93 and 0.74 out of 1.00, respectively, with the Kappa scores 0.81 and 0.59. This indicates that annotators reached near-perfect agreement on the relevance of the generated states to the original while showing moderate agreement on the recognizable variation of those states. Besides, the first five data points mentioned in the guidelines were used as a sanity check to ensure that the annotators correctly understood and applied the evaluation criteria. This ensures the overall quality of the generated states.
5.3 Case Study
Text Image
Query [grounding] After I find 1 transparent mailbox front 2x2. Tell me what to do for the current situation. [Uncaptioned image]
Reference Insert the two tabs at the bottom of the mailbox casing into the two small notches on the mailbox front.
mPLUG-OWL2 Place the mailbox front 2x2 next to the block on the table.
Fine-tuned Put it on the rightmost edge of the high part of the structure.
BLIP2 Put the mailbox front 2x2 horizontally on the table, clasp to the back. Put the mailbox back 2x2 vertically on top, slope to the left.
Fine-tuned Place the mailbox front 2x2 on the left side of the structure to complete the assembly.
LlaVA In the image, there is a small Lego-like object with a red liquid inside of it. The object is placed in a blue sky, and it appears to be a 3D model. To create a transparent mailbox, you need to find a transparent mailbox front that is 2x2 in size. Once you find it, you can use it as the base for your transparent mailbox. You can then place the small Lego-like object inside the mailbox to complete the scene.
Fine-tuned Put it on the mailbox.
QWEN-VL Put the transparent mailbox front 2x2 on the right side of the white arch.
Fine-tuned Put it on the table.
InstructBlip Place it horizontally on the table, clasp to the back.
Fine-tuned Put the mailbox front on the left side of the mailbox base.
MiniGPT-v2 Sure! Here’s what you need to do with the pieces you have collected: 1. Place the transparent mailbox front 2x2 on top of the plate 1x2, aligned with the short side of the plate. Now you have completed the current situation. Let me know if you need help with anything else!
Fine-tuned Put the mailbox front 2x2 horizontally in front of the previous piece.
Otter In view of the reasoning ‘A mailbox is small and square with a see through window.’, the optimal answer should be: build structure.
MiniGPT-4 Attach the transparent mailbox front to the base structure, aligning the tabs with the notches.
GPT-4o Ensure the transparent mailbox front 2x2 is securely attached by inserting the small shaft into the round brick 1x1 before proceeding with the next step.
Table 3: An example of scene understanding results. Words highlighted in orange represent manipulation verbs, red and blue indicate noun entities and their adjectives, and purple denotes positions. The word with a background color orange indicates the task-specific token and the words with a background color green is obtained from the previous step’s textual instruction manual.
In this section, we present illustrative examples to highlight the capabilities and challenges in fine-grained vision-language understanding, particularly in achieving overlap with the reference and accurately recognizing entities and their properties, such as size, color, and etc.
Table 3 present an example showcasing the intuitive results of the prevailing VLMs on scene understanding. We obtain the following observations: First, most models are able to correctly identify the core theme entity “mailbox front” (highlighted in green), with the exception of Otter and GPT-4o when performing scene understanding. However, generating a complete set of properties (e.g., “transparent” highlighted in blue and size “2x2” in pink) to describe these entities accurately and without ambiguity remains a significant challenge. Second, providing the correct assembly operations with precise positional instructions is highly challenging. Although LlaVA and MiniGPT-v2 generate informative instructions, they are not factually accurate in terms of the assembly process.
6 Related Work
6.1 Vision-Language Datasets in AR
Dataset COIN HowTo100M TEACh MIMIC-IT HoloAssist EPIC-KITCHENS Assembly101 Cross-task RareAct LEGO-VLM (Ours)
T1
T2
T3
Size 11,827 23,611 3,215 2.8M 350 89,977 4,321 4,713 7,607 35,612
Table 4: Comparison of datasets for assembly tasks towards AR training regarding the capabilities of tasks T1, T2, and T3. The symbols ✓and ✗indicate the presence or absence of each capability.
Various datasets have been developed to enhance AR assembly tasks by supporting key capabilities T1, T2, and T3. A comparative summary is presented in Table 4.
COIN provides a strong foundation for sequential task analysis with comprehensive annotations, aiding research in multimodal learning (Tang et al., 2019). HoloAssist captures egocentric human-AI interactions using mixed-reality headsets, offering valuable real-world insights (Wang et al., 2023b). HowTo100M, with its extensive collection of 136 million video clips and transcribed narrations, is highly effective for text-to-video retrieval but lacks fine-grained object annotations (Miech et al., 2019). TEACh focuses on interactive dialogues in domestic environments, improving dialogue modeling but not prioritizing detailed object detection (Padmakumar et al., 2021).
While some datasets excel in action recognition, they often lack the procedural depth necessary for effective training. Assembly101, with over 4,000 toy assembly videos, does not include step-by-step instructional details (Sener et al., 2022). CrossTask facilitates weakly supervised learning by leveraging narrations and step lists but lacks temporal annotations to clarify action sequences (Zhukov et al., 2019). EPIC-KITCHENS provides extensive annotations on unscripted kitchen activities, yet it does not offer structured procedural guidance (Damen et al., 2018). RareAct presents unique interactions, challenging models to interpret complex actions without explicit instructions (Miech et al., 2020).
Stanescu et al. (2023a) introduce a state-aware prior that significantly improves object detection in assembly tasks like furniture and Lego construction. However, it lacks textual descriptions of object states and does not address fine-grained detection (Stanescu et al., 2023b). Meanwhile, MIMIC-IT contributes 2.8 million multimodal instruction-response pairs, enriching conversational modeling but lacking scenario-specific descriptions that would enhance task guidance (Li et al., 2023a).
To bridge these gaps, we conduct a comparative analysis evaluating dataset coverage of essential components—objects, states, and scenes—critical for developing AR-guided assistants. Additionally, we introduce a new simulated dataset featuring nearly 400 objects, systematically alternating between object detection and assembly functions. This dataset is designed to optimize the training and evaluation of advanced VLMs, improving their ability to handle fine-grained assembly tasks.
6.2 Towards Fine-grained VLMs
Recent advancements in vision-language models (VLMs) have broadened their applications, enhancing accessibility, search indexing, and interactive content retrieval (Fan et al., 2024; Zhang et al., 2024; Du et al., 2022; Gan et al., 2022). These models generate captions, describe images (Zhou et al., 2020; Hu et al., 2022), facilitate visual QA (Bazi et al., 2023). For instance, they can accurately identify objects in images based on phrases like “the red car” (Subramanian et al., 2022). Furthermore, these models facilitate conversational interactions about visual elements (Chen et al., 2022), which makes virtual assistants more adept at handling inquiries related to images and videos, which makes virtual assistants more adept at handling inquiries related to images and videos. Additionally, VLMs can support multilingual content (Gwinnup & Duh, 2023), benefiting assistive technologies and education (Chi et al., 2020; Wang et al., 2024a).
Despite progress, VLMs often lack fine-grained modeling for scene understanding, object recognition, and error detection. While CLIP demonstrates strong generalization (Radford et al., 2021), models like EfficientVLM (Wang et al., 2023a), MiniGPT-v2 (Chen et al., 2023), and Qwen-VL (Bai et al., 2023) struggle with intricate tasks. OSCAR (Li et al., 2020) and VisionLLM (Wang et al., 2024b) improve image-text alignment but face challenges in context-specific adaptation. The Otter model advances sequential task management yet falls short in detailed analysis (Li et al., 2023b).
Efforts in procedural video representation (Zhong et al., 2023) and instructional task graphs (Ashutosh et al., 2024) highlight potential VLM integration but reveal gaps in error correction. To address these challenges, we propose a vision-language architecture optimized for fine-grained tasks, advancing training and benchmarking of VLMs.
7 Conclusion
Vision-language models (VLMs) have significantly advanced in recent years, enabling AI systems to interpret and generate textual descriptions of visual content. However, despite their success in general vision-language tasks, these models often struggle with fine-grained understanding, particularly in structured instructional scenarios. This research explores fine-grained vision-language modeling for manual-guide AR training, focusing on scene understanding, object detection, and state detection—areas where VLMs often struggle. We developed a specialized dataset from LEGO instruction manuals, designing fine-grained tasks to evaluate VLM performance in tracking assembly sequences and interpreting instructions. Through a one-for-all architecture, we assessed existing models and found them lacking in precision for instructional tasks, highlighting the need for improvement. Future work will analyze model failures and develop enhanced VLMs with better accuracy and contextual understanding. Beyond technical advancements, our research aims to empower blind and visually impaired individuals by enabling AI-driven learning tools, promoting greater accessibility and independence.
Appendix A Analysis
A.1 Impact of Query Construction
We explore the impact of query construction on the QWEN-VL model in the early phase to identify the optimal configuration for achieving the best inference results.
Setting T1 T2 T3
Instruction Example T_ACC $\uparrow$ ROUGE $\uparrow$ BLEU $\uparrow$ E_ACC $\uparrow$ IOU $\uparrow$ F1-State $\uparrow$ FPR $\downarrow$
11.90 29.57 5.19 32.41 25.60 39.77 97.99
6.84 24.63 4.58 5.56 22.01 34.00 97.50
10.43 23.00 4.24 26.37 15.55 34.50 97.20
5.27 11.02 1.30 3.70 0.10 25.50 98.40
Table 5: A study of query construction results across various tasks and metrics. The symbol $\uparrow$ indicates higher is better, $\downarrow$ indicates lower is better, and values in bold represent the highest results performance.
To this end, we conducted experiments on various query settings to determine which components most significantly influence performance. The model was chosen for this study due to its strong results under our base setting, which included both instruction directives and in-context examples.
As shown in Table 5, our findings highlight the critical role of query design across different tasks. The results indicate that providing both instructions and examples yields the best overall performance, demonstrating their complementary roles in guiding inference.
In the scene understanding task (T1), the presence of instructions plays a critical role in improving accuracy. When instructions are included, theme accuracy (T_ACC) increases significantly, suggesting that clear guidance helps the model better understand scene context and thematic elements. Furthermore, both ROUGE and BLEU scores, which evaluate the structure and coherence of generated outputs, improve when at least one of the two components (instructions or examples) is present. The lowest performance in this task occurs when neither component is provided, confirming that structured input is essential for meaningful inference.
In the object detection task (T2), the results suggest that examples contribute more to entity recognition and output formatting than instructions alone. The entity accuracy (E_ACC) drops sharply from 32.41% to 5.56% when examples are removed, even when instructions remain. Similarly, IoU, which measures how well the output adheres to the expected format, significantly decreases in the absence of examples. This highlights that examples provide concrete reference points, enabling the model to generate well-structured and accurate entity labels.
In the state detection task (T3), despite improvements in other metrics with enhanced query settings, the false positive rate (FPR) remains consistently high (above 97%) across all conditions. This suggests that the model has a strong affirmative bias, likely defaulting to generating positive responses regardless of query refinements. While instructions and examples improve structured outputs, they do not sufficiently mitigate the issue of excessive false positives.
These findings emphasize the importance of query design in optimizing model performance. Combining instructions with instruction directives and in-context examples is crucial for achieving high accuracy, structured outputs, and reliable entity recognition. However, the persistent issue of high FPR suggests the need for further refinements, such as introducing negative examples, confidence calibration, or additional fine-tuning strategies to reduce overconfidence in affirmative responses. Future work should explore these approaches to enhance the model’s robustness and reliability in diverse inference tasks.
A.2 Case Study of Object Detection
Text Image
Query [object] Please identify and highlight the positions of 1 light purple brick 2x6. Providing the positions in the format: <p>object</p> <Xleft><Ytop><Xright><Ybottom>, with X and Y coordinates normalized to [0,100]. <Xleft> and <Ytop> for the top-left corner. <Xright> and <Ybottom> for the bottom-right corner. [Uncaptioned image]
Reference <p>1 light purple brick</p> <58><15><92><35>
mPLUG-OWL2 <p>1 light purple brick</p> <125><48><45><97>
Fine-tuned <p>1 light purple brick</p> <60><24><69><35>
BLIP2 <p>1 light purple brick</p> <57><15><82><35><p>and<p>1 dark purple brick 2x6</p> <82><25><93><36>
Fine-tuned <p>1 light purple brick</p> <58><15><92><35>
LLaVA The light purple brick is located at the center of the image, surrounded by yellow and black bricks.
Fine-tuned <p>1 light purple brick</p> <62><15><87><35>
QWEN-VL <20><61><30><78>
Fine-tuned <p>1 light purple brick</p> <57><15><82><35>
InstructiBLIP <p>1 light purple brick 2x6</p> <57><15><82><35>
Fine-tuned <p>1 light purple brick</p> <58><15><92><35>
MiniGPT-v2 <p>1 light purple brick</p> <62><20><88><40>
Fine-tuned <p>1 light purple brick</p> <57><17><85><37>
Otter The 1 light purple brick 2x6 is on the left side of the image.
MiniGPT-4 Sure! Here are the positions of the light purple brick in the image: Xleft = 0, Ytop = 25, Xright = 30, Ybottom = 50.
GPT-4o from PIL import Image # Load the image image_path $\dots$
Table 6: An example of object detection results. The object is highlighted in light purple, and its position is highlighted in blue.
Table 6 present an example of object detection task. We obtain the following observations: First, only InstructiBLIP and Otter generate the full description of the object “1 light purple brick 2x6” without disambiguation. Second, none of the models accurately predict the precise positions as outlined in the reference.
Appendix B Enlightening Discussion
B.1 Limitations
The quality of object and state detection in our dataset heavily depends on the performance of current VLMs. While we have assessed the data quality, it is important to note that this is a synthetic dataset with only limited human intervention for corrections. In the object detection task, the model generates coordinates used to highlight detected objects with bounding boxes, though these coordinates are not meant for direct user interaction. In contrast, the outputs of the state detection and scene understanding tasks are text-based and designed to be user-friendly. However, we have not yet accounted for 3D scenes or objects. We evaluated a few existing VLMs as benchmarks, but their performance remains suboptimal. Future work will focus on improving fine-grained VLMs to enhance their effectiveness in these tasks.
B.2 Broader Impact
Our work has significant implications for accessibility and inclusive learning, particularly in assisting blind and visually impaired users (Gadiraju et al., 2020; Hao et al., 2024; Karamolegkou et al., 2024; Zhao et al., 2024). By integrating VLMs into conversational training assistants, we aim to create equitable learning opportunities in AR-based training environments. This could enable individuals with visual impairments to engage in hands-on assembly tasks through AI-powered guidance, reducing barriers to education and skill development.
Beyond accessibility, our research contributes to industrial applications, particularly in manufacturing, robotics, and assembly training. Fine-grained vision-language reasoning can enhance AR-based instruction systems, reducing errors, improving efficiency, and enabling more intuitive human-AI collaboration in high-stakes environments like car manufacturing and industrial product assembly. By bridging the gap between VLMs and real-world AR applications, our work also fosters more intelligent and interactive training experiences, benefiting industries that rely on precise, step-by-step assembly guidance.
B.3 Future Work
A key area for future research is improving VLMs for fine-grained spatial reasoning. Current models struggle with detailed object and state recognition, limiting their ability to provide accurate real-time feedback in AR training. Advancements in pretraining strategies, multimodal fusion techniques, and architectural design could enhance spatial understanding in AR-based assembly tasks.
Expanding the dataset is another critical direction. While our dataset focuses on LEGO-based assembly tasks as a simplified simulation, future work should explore real-world industrial assembly scenarios, such as furniture construction, electronic device assembly, and mechanical repairs. This would improve model generalization and practical applicability in professional training settings.
Additionally, integrating real-time multimodal interaction is essential for developing dynamic AR assistants. Future research should explore VLMs that can process live user input, track assembly progress, and provide adaptive feedback based on visual and textual cues. This would move beyond static dataset-based evaluation toward fully interactive AR guidance systems.
Finally, trust and explainability in human-AI collaboration remain crucial. Understanding how users interact with AI-powered AR assistants can help improve trust, usability, and transparency. Future work should investigate methods for making VLMs more interpretable and reliable, ensuring that AI-generated instructions align with human expectations and real-world assembly requirements.
Appendix C Implementation Details
C.1 Implementational setup
During data synthesis, we utilized MiniGPT4-v2777https://github.com/Vision-CAIR/MiniGPT-4 as the VLM, employing straightforward rules and human correction to create vision-language data. To evaluate the performance of benchmark models, we randomly sampled 25% of the dataset to serve as the test set. All inference and fine-tuning experiments were conducted on NVIDIA A100 SXM4 40GB GPUs. We set the learning rate to 1e-5 and the batch size to 16 for our experiments. The AdamW optimizer (Loshchilov & Hutter, 2017) was employed to facilitate efficient training, which spanned 3 epochs.
C.2 Query Templates
In this paper, we investigate three fine-grained vision-language tasks and have devised specific query templates for each, as demonstrated in Table 7.
T1: Scene Understanding
Token: [grounding]
Selection Rule: <step_symbol> with <step>. Instruction from this step.
Format-Directive: Tell me what to do for the current situation.
T2: Object Detection
Token: [object]
Selection Rule: <step_symbol> with <eop>, Word starts with “find” or “collect”. Instruction from this step.
Format-Directive: Providing the positions in the format: <p>object</p> {<|Xleft><Ytop><Xright><Ybottom>}, with X and Y coordinates normalized to [0,100]. <Xleft> and <Ytop> for the top-left corner. <Xright> and <Ybottom> for the bottom-right corner.
T3: State Detection
Token: [state]
Selection Rule: <step_symbol> with <step>. Previous <step_symbol> is <eop>. Instruction from this step and the previous step.
Format-Directive: Just tell me Yes or No.
Table 7: Query Templates for Fine-Grained VLM Tasks.
C.3 Dataset Details
C.3.1 Data Schema in JSON Format
This section presents the JSON data schema for integrating multimodal data (visual and textual) and fine-tuning VLM. The schema defines key components that structure and organize data efficiently, ensuring seamless preprocessing before final compilation.
  "instruction_id": 95,
  "text": [ "Find 2 transparent red flat tiles 1x2." ],
  "VLM": {
    "img_path": "...",
    "parts_img_path": "None",
    "id_sign": "Car",
    "step_num": 28,
    "step_class": "eop",
    "other_sign": "AG",
    "background_rect": "None",
    "color_list": [],
    "bound_color": [],
    "task_label": "object",
    "query": "Find 2 transparent red flat tiles 1x2.",
    "MiniGPTv2_output": "<p>2 transparent red flat tiles</p> {<28><17><73><50>}"
  "img": "...",
  "entities": [
    { "this_line": "Find 2 transparent red flat tiles 1x2." },
    { "verb": "Find" },
    { "det": "2" },
    { "colorinfo": "transparent red" },
    { "theme": "flat tiles 1x2" },
    { "suffix": "." } ]
Table 8: An example of a JSON schema for integrated multimodal data.
The key fields related to VLM in the JSON schema include:
  • img_path: Stores the local path of images downloaded via web scraping, pointing to the original files fetched from the web.
  • id_sign: Extracted from the image filename, this field identifies the sub-assembly session (e.g., “CAR” in this example).
  • step_num: Represents the sequence number of the assembly step (e.g., step 28), tracking the assembly’s progression.
  • step_class: A classification tag (e.g., “eop” for “end of part”) indicating an object-finding step in the assembly process.
  • MiniGPTv2_output: Stores processed output from MiniGPTv2, containing annotations or modifications essential for subsequent dataset creation.
  • parts_img_path, bound_color, color_list: Reserved for assembly steps classified as “step” and crucial for state detection dataset generation:
    • color_list: Records all possible boundary colors used throughout the assembly session.
    • bound_color: Consolidates the most probable boundary colors through a voting mechanism, ensuring consistency across the dataset.
To support fine-tuning, our default JSON format follows the Llava model’s fine-tuning data structure, ensuring an optimized and adaptable framework for general use cases. An example of the JSON schema for VLM fine-tuning is illustrated in Table 9.
  "id": "lego-60263-ocean-mini-submarine-readscr-12",
  "image": "sm01-island-0_step_Step%20%23%23E.png",
  "conversations": [
      "from": "human",
      "value": "[grounding] After I find 1 reddish brown chest bottom 2x4.
      Tell me what to do for the current situation."
      "from": "gpt",
      "value": "Put it horizontally on the table."
Table 9: An example of a JSON schema for VLM fine-tuning.
The full dataset, comprising JSON and PNG files, occupies approximately 11.8 GB. For specialized requirements, such as those for the Otter model, we convert all text and image data into Parquet files.
C.3.2 An Example of LEGO Instruction Manual
Figure 5 shows an example of a part of the instruction manual we crawled from the official LEGO website.
Figure 5: An example of LEGO instruction manual. It consists of a summary section at the beginning, followed by three sequential instruction steps. Each step includes textual instructions paired with corresponding images to guide the assembly process.
The LEGO audio instruction manual is a structured guide designed to assist blind and visually impaired users in assembling LEGO sets accurately. The key sections of a LEGO instruction manual include: (1) a summary providing general information, such as topics and candidate parts for assembly, followed by (2) a sequence of multimodal step instructions. Each step consists of textual instructions paired with an illustrative image, with key functional phrases, such as theme entities, highlighted for clarity provided by LEGO. With an average of 215.3 steps per manual, we segmented each manual into multiple assembly sessions, each consisting of a paired sequence: an object-seeking step followed by an object-assembly step.
C.3.3 Illustration of Data Creation.
Figure 6 illustrates the process of creating the dataset for task T2. It begins by iterating through object-seeking steps, where users locate specific objects. Next, images are processed to generate the object detection dataset, consisting of query-image-text triplets. Queries are formed using a template with the task token "[object]" and the current instruction. Responses include positional text indicating object coordinates, initially inferred using MiniGPT-v2. A composite image is then created by merging the target image with three randomly sampled images, and the coordinates are adjusted accordingly.
Figure 6: An illustration of object detection data creation.
Figure 7 illustrates the process of generating the dataset for task T3. The procedure begins by iterating through object-assembly steps, pairing each with its preceding step where users identified parts to assemble. Next, part segmentation is performed by detecting boundary colors using K-means clustering and Hue filtering, followed by manual correction. Boundaries are highlighted in distinct colors to mark assembly objects. Bounding boxes are then added to segmented objects, and natural perturbations are introduced by shifting assembled parts within the background box. Each correct assembly state serves as a positive sample, while three perturbed states act as negative samples. Finally, the dataset is constructed, where queries integrate task tokens, current and previous instructions, and responses indicate correct or incorrect assembly states.
Figure 7: An illustration of synthesized negative samples for state detection.
C.3.4 Examples in Data Quality Assessment
Figure 8 and 9 illustrate examples of the quality assessment for generated data in object detection (T2) and state detection (T3), respectively. We randomly selected 100 data examples per task to ensure a fair assessment. Three annotators were provided with comprehensive guidelines on how to evaluate each dataset based on multiple criteria, accompanied by relevant examples. The first five data points mentioned in the guidelines were used as a sanity check to ensure that the annotators correctly understood and applied the evaluation criteria. If annotators fail the sanity check, they undergo retraining until they pass. We evaluate annotations using the average scores of four criteria (i.e., entity disambiguity, boundary precision, state relevance, and state identifiability) and calculate Kappa values to assess annotation agreement.
Figure 8: An example of evaluating data quality of task T2.
Figure 9: An example of evaluating data quality of task T2.

### v2

@@
-Fine-Grained Vision-Language Modeling for Multimodal Training Assistants in Augmented Reality
+LEGO Co-builder: Exploring Fine-Grained Vision-Language Modeling for Multimodal LEGO Assembly Assistants
 Abstract
+Vision-language models (VLMs) are facing the challenges of understanding and following multimodal assembly instructions, particularly when fine-grained spatial reasoning and precise object state detection are required.
+In this work, we explore LEGO Co-builder, a hybrid benchmark combining real-world LEGO assembly logic with programmatically generated multimodal scenes.
+The dataset captures stepwise visual states and procedural instructions, allowing controlled evaluation of instruction-following, object detection, and state detection.
+We introduce a unified framework and assess leading VLMs such as GPT-4o, Gemini, and Qwen-VL, under zero-shot and fine-tuned settings.
+Our results reveal that even advanced models like GPT-4o struggle with fine-grained assembly tasks, with a maximum F1 score of just 40.54% on state detection, highlighting gaps in fine-grained visual understanding.
+We release the benchmark, codebase, and generation pipeline to support future research on multimodal assembly assistants grounded in real-world workflows.
+Introduction
+Multimodal instruction-following assistants are gaining increasing relevance in domains requiring precise procedural understanding, such as furniture construction (You et al.
+2022), automotive manufacturing (Bellalouna et al.
+2020), and industrial product assembly (Funk et al.
+2017).
+These tasks demand step-by-step reasoning, spatial awareness, and accurate interpretation of visual and textual instructions—capabilities that current AI systems still struggle to reliably deliver.
+Figure 1: An example of incorrect state detection by GPT-4o during LEGO assembly.
+Given an image and text instructions, the model fails to accurately recognize an incorrectly placed LEGO part (highlighted in green), demonstrating the challenge of fine-grained vision-language alignment.
+Figure 2: An illustration of the proposed three VLM tasks, highlighting the core capabilities required by the one-for-all VLM architecture for LEGO assembly based on the procedural instruction manual.
+Traditional computer vision research has tackled various purely visual tasks, including action segmentation, recognition, anticipation, and object state detection (Damen et al.
+2018; Miech et al.
+2020; Wang et al.
+2023b; Gao et al.
+2024).
+Procedural activity understanding has also been explored using video datasets with or without textual narration (Tang et al.
+2019; Zhukov et al.
+2019; Li et al.
+2023a; Sener et al.
+2022).
+Some methods incorporate textual alignment, such as narration-video grounding or instruction-conditioned generation (Padmakumar et al.
+2021; Miech et al.
+2019), but often rely on separate architectures for each task type.
 […]
-VLM are essential for enabling AI-powered smart assistants to interpret and reason in multimodal environments.
-However, their application in augmented reality (AR) training remains largely unexplored.
-In this work, we introduce a comprehensive dataset tailored for AR training, featuring systematized vision-language tasks, and evaluate nine state-of-the-art VLMs on it.
-Our results reveal that even advanced models, including GPT-4o, struggle with fine-grained assembly tasks, achieving a maximum F1 score of just 40.54% on state detection.
-These findings highlight the demand for enhanced datasets, benchmarks, and further research to improve fine-grained vision-language alignment.
-Beyond technical contributions, our work has broader social implications, particularly in empowering blind and visually impaired users with equitable access to AI-driven learning opportunities.
-We provide all related resources, including the dataset, source code, and evaluation results, to support the research community.
-1 Introduction
-Training assistants are in high demand in industrial assembly tasks, including furniture production (You et al., 2022), car manufacturing (Bellalouna et al., 2020), and industrial product assembly (Funk et al., 2017).
-As these tasks often require high levels of precision and cognitive effort, augmented reality (AR) has emerged as a promising technology to support workers, providing a multimodal, interactive, user-centric environment (Hou & Wang, 2013; Botto et al., 2020; Dalim et al., 2020).
-AR training systems have been extensively studied for their ability to enhance user experience and task efficiency (Büttner et al., 2020; Blattgerste et al., 2021).
-However, most existing research focuses on interaction design and usability studies, rather than the development of intelligent, instruction-following assistants that can understand and process complex manual-based assembly tasks in AR environments.
-[…]
-VLM offer a powerful solution for bridging visual and textual modalities by effectively interpreting and integrating visual inputs with corresponding textual instructions.
-Traditional computer vision studies have explored various purely visual tasks, such as action segmentation, recognition, anticipation, and object and state detection (Damen et al., 2018; Miech et al., 2020; Wang et al., 2023b; Gao et al., 2024).
-Some studies focus on analyzing procedural activities in videos without fixed instructions (Tang et al., 2019; Sener et al., 2022), while others incorporate textual instructions to enhance comprehension.
-These approaches include using instructional queries for response generation(Li et al., 2023a; Padmakumar et al., 2021), narration-video alignment(Zhukov et al., 2019), and text-to-video retrieval(Miech et al., 2019).
-However, none of these studies explicitly address instruction-following in manual assembly tasks, where models must accurately interpret step-by-step procedural instructions and track fine-grained object states within a structured pipeline.
-The recent advancements in large language models and vision-language models have made multimodal reasoning more accessible, with models such as GPT-4(Achiam et al., 2023), LLaVa (Liu et al., 2023), QWEN-VL (Bai et al., 2023), BLIP2 (Li et al., 2023c), and MiniGPT-v2 (Chen et al., 2023).
-The core idea is to excel in multimodal tasks by integrating visual encoders with fundamental LLMs.
-Some early-stage explorations have attempted to apply VLMs in AR to align visual contexts with language representations to showcase the capabilities of LLM agents (Pei et al., 2024).
-However, these efforts have primarily focused on coarse-grained visual features, such as detecting objects and their coordinates, rather than understanding complex procedural assembly tasks in detail.
-A key challenge in AR-based assembly training is the fine-grained interpretation of object states and spatial relationships.
-Figure 1 illustrates this limitation: when prompted with an image and text instructions for LEGO assembly, GPT-4o fails to correctly identify an incorrect assembly state, where a newly added LEGO part (in green) is misplaced.
-This example highlights a crucial weakness – current VLMs struggle with precise state detection and fine-grained assembly verification, making them unreliable for industrial AR training scenarios.
-Figure 1: An example of incorrect state detection by GPT-4o during LEGO assembly.
-Given an image and text instructions, the model fails to accurately recognize an incorrectly placed LEGO part (highlighted in green), demonstrating the challenge of fine-grained vision-language alignment in assembly tasks.
-To address these challenges, we introduce a comprehensive vision-language benchmark and dataset designed for AR-based assembly training.
-Specifically, we chose LEGO brick assembly in AR as a simplified simulation and created a synthetic multimodal dataset to address vision-language challenges, i.e., complex scene understanding, object detection, and state detection.
-We developed a unified, fine-grained VLM architecture capable of handling a broad range of vision-language tasks in AR and validated the feasibility of nine prevailing models as benchmarks.
-Notably, this research is designed for various manual-guided assembly assistance in a multimodal environment, though we specifically demonstrated it using LEGO assembly.
-We list contributions below: (1) Developing a unified VLMs-based architecture to benchmark fine-grained instruction-following capabilities in AR training.
-(2) Evaluating nine prevailing VLMs, both without and with fine-tuning, on a synthetic multimodal dataset tailored for AR-based LEGO assembly.
-(3) Creating and releasing a dataset containing systematic vision-language tasks for AR training, along with a method for synthesizing high-quality multimodal data.
-2 Fine-grained Vision-Language Modeling
-2.1 Task definition
-As illustrated in Figure 2, we investigate manual-guided LEGO assembly in AR and define the following three vision-language tasks: (T1) Scene Understanding.
+VLM offer a promising solution by bridging visual and textual modalities, enabling models to align procedural language with visual observations.
+Recent advances in large-scale models such as GPT-4o (Achiam et al.
+2023), LLaVa (Liu et al.
+2023), Qwen-VL (Bai et al.
+2023), BLIP-2 (Li et al.
+2023c), MiniGPT-v2 (Chen et al.
+2023), and more recent reasoning-augmented variants like VLM-R1 (Shen et al.
+2025), GLM-4.1v-thinking (Hong et al.
+2025), and Gemini2.5 (Comanici et al.
+2025), have demonstrated strong performance on a range of general-purpose multimodal tasks.
+However, most existing benchmarks focus on coarse-grained capabilities such as object recognition, captioning, and visual question answering, and do not evaluate the detailed, procedural understanding required for step-by-step manual assembly.
+While recent VLMs research (Laurençon et al.
+2024; Cheng et al.
+2024; Chen et al.
+2024a) aim to unify vision-language reasoning under a single architecture, challenges remain—particularly for tasks that require dense and fine-grained perception (Wei et al.
+2024; Rahmanzadehgervi et al.
+2024; Chen et al.
+2024b).
+In this work, we address the need for instruction-following benchmarks that evaluate a model’s ability to reason over both visual and linguistic input in detail.
+We focus on LEGO brick assembly as a representative manual task that is richly structured, visually complex, and procedurally grounded.
+Figure 1 illustrates the limitations of current models in such tasks.
+When prompted with a visual scene and corresponding instruction, GPT-4o fails to detect a misplaced LEGO piece, highlighting the challenge of fine-grained visual scene understanding and object state recognition.
+This motivates the need for benchmarks that go beyond high-level visual understanding and evaluate precise spatial compliance with procedural steps.
+To address this gap, we present LEGO Co-builder, a benchmark and dataset designed to evaluate fine-grained instruction-following capabilities in LEGO assembly tasks.
+Our dataset includes multimodal sequences consisting of visual snapshots, object states, and procedural steps derived from human-crafted manuals.
+We introduce a unified task formulation and evaluate nine leading VLMs in both zero-shot and fine-tuned settings.
+Our contributions are summarized as follows: (1) We develop a unified vision-language architecture to benchmark fine-grained, multimodal, instruction-following capabilities in procedural manual-guided assembly tasks.
+(2) We evaluate prevailing VLMs on a dataset of LEGO assembly sequences with grounded visual and textual supervision, under both zero-shot and fine-tuned settings.
+(3) We release the dataset, benchmarking code, and a modular synthetic data generation pipeline to support future research in multimodal instruction-following.
+Societal Impact.
+This work advances the development of multimodal AI assistants that can transform how people learn and perform complex physical tasks.
+By enabling fine-grained vision-language understanding, our benchmark supports the creation of multimodal educational tools that combine visual input, language, and potentially augmented reality (AR) to enhance hands-on learning experiences.
+These systems can improve training in domains like education and industrial training.
+Moreover, the ability to interpret and verify procedural steps visually opens up promising avenues for assistive technologies—particularly for blind or visually impaired learners—by providing real-time, AI-driven guidance through tasks that were previously inaccessible.
+In this way, our research supports both innovation in teaching modalities and greater inclusivity in skill development.
+Fine-Grained Vision-Language Modeling
+Task Definition
+We investigate manual-guided LEGO assembly and define the following three vision-language tasks (See in Figure 2):
+  1. (T1)
+    Scene Understanding.
 Given the current step’s image $V_{t}$ and the task-specific query $Q_{t}$, the model outputs a response containing the scene description for the current assembly step $I_{t}$.
@@
 It assesses the model’s ability to accurately describe the current assembly step from an instruction manual, including but not limited to generating object representations, their properties, and the assembly procedure.
-(T2) Object Detection.
-Given a seeking step’s image $V_{t^{\prime}}$ ($t^{\prime}=t-1$) and task-specific query $Q_{t^{\prime}}$, the model outputs a response containing the positional text $P_{t^{\prime}}$, which includes identified object and its coordinate formatted as “<Object> <Xleft> <Ytop> <Xright> <Ybottom>”.
+  2. (T2)
+    Object Detection.
+Given a seeking step’s image $V_{t^{\prime}}$ ($t^{\prime}=t-1$) and task-specific query $Q_{t^{\prime}}$, the model outputs a response containing the positional text $P_{t^{\prime}}$, which includes identified object and its coordinate formatted as “[Object] [Xleft] [Ytop] [Xright] [Ybottom]”.
 The query $Q_{t^{\prime}}$ is the task description contextualized with the corresponding textual instruction $I_{t^{\prime}}$.
 It assesses the model’s ability to accurately identify and locate objects in the scene, ensuring they are positioned correctly for the next assembly action.
-(T3) State Detection.
+  3. (T3)
+    State Detection.
 Given an assembly step’s image, either the original manual image $V_{t}$ or a generated negative sample $\tilde{V}_{t}$, and task-specific query $Q_{t}$, the model outputs a response indicating a correct or incorrect state.
 It evaluates the model’s ability to verify the accuracy of the assembly progression, determining if the assembly action has been correctly completed.
-These tasks have been crafted to rigorously test and demonstrate the capabilities of VLMs in an AR training, focusing on their ability to understand and interpret complex multimodal inputs.
-Figure 2: An illustration of the proposed three VLM tasks, highlighting the core capabilities required by the one-for-all VLM architecture for LEGO assembly based on the procedural instruction manual.
-2.2 One-for-all architecture
+These tasks have been crafted to rigorously test and demonstrate the capabilities of VLMs in LEGO Assembly, focusing on their ability to understand and interpret complex multimodal inputs.
+One-for-all Vision-Language Architecture
 We investigated existing VLMs and derived a universal architecture for the proposed tasks, as illustrated in Figure 3.
@@
 Figure 3: The architecture for one-for-all vision-language modeling.
-It integrates a vision encoder, a large language model, and a vision-language projector as core functional modules, along with a task-specific query for task adaptation.
-3 Dataset Creation
-3.1 Manual Crawling and Scene Data Matching
-This subsection outlines the procedure for creating the dataset for task T1.
-First, we collected 65 instruction manuals from the official LEGO website [^1] https://legoaudioinstructions.com/instructions, designed to help blind and visually impaired users assemble LEGO sets accurately.
+It integrates a vision encoder, a large language model, and a vision-language projector as core functional modules, with a task-specific query for task adaptation.
+Dataset Creation
+Manual Crawling and Scene Data Matching
+We outline the procedure for creating the dataset for task T1.
+First, we collected 65 official LEGO instruction manuals (LEGO 2024), designed to help blind and visually impaired users assemble LEGO sets accurately.
 Each instruction manual comprises several elements: (1) step-by-step textual instructions; (2) corresponding image for each step; and (3) tags provided by LEGO include instruction and image tags.
@@
 Next, we iterated through all sessions to generate the scene understanding dataset, $D_{T1}=\{(Q_{t},V_{t},A_{t})\}|_{t=0}^{|D_{T1}|}$, where each element is a triplet of (query, image, text), and the query $Q_{t}$ is constructed by filling the query template with the task token “[grounding]” and the previous instruction $I_{t-1}$ as context.
-3.2 Object Position Inference
-This subsection outlines the procedure for creating the dataset for task T2, as illustrated in Figure 6.
+Object Position Inference
+This subsection outlines the procedure for creating the dataset for task T2.
 First, we iterated through all object-seeking steps, where users are asked to find specific objects.
@@
 The query $Q_{t^{\prime}}$ is constructed by filling the query template with the task token “[object]” and the current instruction $I_{t^{\prime}}$.
-The response $A_{t^{\prime}}$ contains a positional text $\tilde{P}_{t^{\prime}}$ formatted as “<Object> <Xleft> <Ytop> <Xright> <Ybottom>”.
-This is initially generated by querying the image using MiniGPT-v2 (Chen et al., 2023), the state-of-the-art model at the time of this work.
+The response $A_{t^{\prime}}$ contains a positional text $\tilde{P}_{t^{\prime}}$ formatted as “[Object] [Xleft] [Ytop] [Xright] [Ybottom]”.
+This is initially generated by querying the image using MiniGPT-v2 (Chen et al.
+2023), the state-of-the-art model at the time of this work.
 The composite image $\tilde{V}_{t^{\prime}}$ is created by combining the current image ${V}_{t^{\prime}}$ with three randomly sampled images.
 Last, the initial coordinates are adjusted to fit the composite image.
-3.3 Variant State Generation
-This subsection outlines the procedure for creating the dataset for task T3, as illustrated in Figure 7.
+Variant State Generation
+This subsection outlines the procedure for creating the dataset for task T3.
 First, we iterated through all object-assembly steps, along with their corresponding previous steps, where users are trained to assemble the parts identified in the prior steps.
@@
 In each assembly session, the objects to be assembled in the current step are tagged in the previous step, with their boundaries highlighted in distinct colors that vary between sessions.
-The highlighted colors are selected based on predominant color detection using K-means clustering [^2] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html, followed by Hue filtering (Chu et al., 2014) and manual correction for accuracy.
+The highlighted colors are selected based on predominant color detection using K-means clustering (Statsmodels 2024), followed by Hue filtering (Chu, Tsuji, and Kato 2014) and manual correction for accuracy.
 Third, we added bounding boxes to the segmented objects.
@@
 The responses $A_{t}$ and $\{\tilde{A}_{t},\tilde{A}_{t}^{{}^{\prime}},\tilde{A}_{t}^{{}^{\prime\prime}}\}$ indicate the correct and incorrect statuses of the assembly state, respectively.
-3.4 Dataset Statistics
+Dataset Statistics
 We summarize the statistics of LEGO-VLM dataset in Table 1.
@@
 Overall, the dataset includes 35,612 vision-language data samples: 5,612 for T1, 19,136 for T2, and 10,864 for T3, respectively.
-Manual Session Scene Object State Step Sample
-Overall Identification Assembly Overall T1 T2 T3
-65 397 5,612 4,784 2,716 10,428 4,814 5,614 35,612 5,612 19,136 10,864
+# Manual # Session # Scene # Object # State
+65 397 5,612 4,784 2,716
+# Step Overall Identification Assembly
+10,428 4,814 5,614
+# Sample Overall T1 T2 T3
+35,612 5,612 19,136 10,864
 Table 1: Statistics of LEGO-VLM dataset.
-4 Experimental Setup
-4.1 Benchmarks
-We benchmark the nine prevailing VLMs: (1) mPLUG-OWL2 (Ye et al., 2024)replaces attention with a modality adapter in a LLM decoder.
-(2) BLIP2 (Li et al., 2023c)uses a two-stage pre-trained Q-Former between an image encoder and a LLM to bridge the vision-language modality gap.
-(3) LLaVa (Liu et al., 2023)combines a visual CLIP encoder and a language decoder Vicuna for general-purpose visual language understanding.
-(4) Qwen-VL (Bai et al., 2023)connects a LLM with a visual encoder using position-aware vision-language adapter towards fine-grained visual understanding.
-(5) InstructBLIP (Dai et al., 2023)explores general-purpose vision-language instruction tuning based on the pretrained BLIP2.
-(6) MiniGPT-v2 (Chen et al., 2023)links a frozen ViT visual encoder with Llama-2 via a projection layer, applicable for diverse tasks via task-specific multimodal instructions.
-(7) Otter (Li et al., 2023a)is tuned on the OpenFlamingo, conditioning the language model on images for multi-modal perception and reasoning.
-(8) MiniGPT-4 (Zhu et al., 2023)integrates a frozen visual ViT&Q-Former encoder and Vicuna via a projection layer, unlocking advanced multimodal capabilities like GPT-4.
-(9) GPT-4o [^3] https://openai.com/index/hello-gpt-4o/is a widely used commercial model accessible via APIs.
-4.2 Evaluation Metrics
+Experimental Setup
+Benchmarks
+We benchmark the following nine prevailing VLMs: (1) mPLUG-OWL2 (Ye et al.
+2024)replaces attention with a modality adapter in a large language model (LLM) decoder.
+(2) BLIP2 (Li et al.
+2023c)uses a two-stage pre-trained Q-Former between an image encoder and a LLM to bridge the vision-language modality gap.
+(3) LLaVa (Liu et al.
+2023)combines a visual CLIP encoder and a language decoder Vicuna for general-purpose visual language understanding.
+(4) Qwen-VL (Bai et al.
+2023)connects a LLM with a visual encoder using position-aware vision-language adapter towards fine-grained visual understanding.
+(5) InstructBLIP (Dai et al.
+2023)explores general-purpose vision-language instruction tuning based on the pretrained BLIP2.
+(6) MiniGPT-v2 (Chen et al.
+2023)links a frozen ViT visual encoder with Llama-2 via a projection layer, applicable for diverse tasks via task-specific multimodal instructions.
+(7) Otter (Li et al.
+2023a)is tuned on the OpenFlamingo, conditioning the language model on images for multi-modal perception and reasoning.
+(8) MiniGPT-4 (Zhu et al.
+2023)integrates a frozen visual ViT&Q-Former encoder and Vicuna via a projection layer, unlocking advanced multimodal capabilities like GPT-4.
+(9) GPT-4o(OpenAI 2024)is a widely used commercial model accessible via APIs.
+(10) Qwen-VL-2.5update Qwen-VL with the latest Qwen LLM (QwenTeam 2024).
+(11) Gemini-2.5-flash(Comanici et al.
+2025)is a thinking model, designed to tackle increasingly complex problems.
+(12) GLM-4.1-thinking(Hong et al.
+2025)is designed to explore the upper limits of reasoning.
+Evaluation Metrics
 […]
-5 Outcomes
-5.1 Benchmark Results
+Outcomes
+Benchmark Results
 We compare the performance of nine prevailing VLMs on the proposed LEGO-VLM dataset, without and with finetuing, as shown in Table 2.
@@
 PEFT (LoRA) w/o w/ w/o w/ w/o w/ w/o w/ w/o w/ w/o w/ w/o w/
-mPLUG-OWL2 Ye et al. (2024) 23.16 29.70 25.23 32.75 3.04 8.46 77.94 93.05 14.23 34.88 36.82 15.00 63.41 35.00
-BLIP2 Li et al. (2023c) 27.85 32.65 29.55 40.35 6.40 12.50 77.35 84.48 34.25 40.57 24.71 28.50 100.00 94.87
-LlaVA Liu et al. (2023) 32.10 35.88 13.63 42.97 0.99 17.73 54.39 72.32 $\emptyset$ 60.98- 25.04 30.00 99.41 50.00
-QWEN-VL Bai et al. (2023) 34.84 37.00 29.57 39.08 5.19 13.13 78.56 89.15 25.60 30.08 39.77 39.53 97.99 96.82
-InstructBLIP Dai et al. (2023) 32.85 32.85- 29.20 36.87 4.91 11.52 79.76 98.16 $\emptyset$ 47.20- 0.00 0.02 1.22 0.00
-MiniGPT-v2 Chen et al. (2023) 33.06 37.52 34.72 32.56 9.81 8.28 84.95 85.91 26.98 25.94 36.76 38.64 60.42 80.72
-Otter Li et al. (2023a) 11.49 12.12 1.28 72.39 $\emptyset$ 35.33 75.88
-MiniGPT-4 Zhu et al. (2023) 34.11 15.05 1.88 87.09 30.20 37.19 67.37
-GPT-4o 44footnotemark: 4 25.81 18.67 2.00 66.67 21.68 40.54 43.06
-Table 2: Benchmarking vision-language models on LEGO-VLM dataset, without (/wo) and with (/w) PEFT using LoRA.
+mPLUG-OWL2 23.16 29.70 25.23 32.75 3.04 8.46 77.94 93.05 14.23 34.88 36.82 15.00 63.41 35.00
+BLIP2 27.85 32.65 29.55 40.35 6.40 12.50 77.35 84.48 34.25 40.57 24.71 28.50 100.00 94.87
+LlaVA 32.10 35.88 13.63 42.97 0.99 17.73 54.39 72.32 $\emptyset$ 60.98- 25.04 30.00 99.41 50.00
+Qwen-VL 34.84 37.00 29.57 39.08 5.19 13.13 78.56 89.15 25.60 30.08 39.77 39.53 97.99 96.82
+InstructBLIP 32.85 32.85- 29.20 36.87 4.91 11.52 79.76 98.16 $\emptyset$ 47.20- 0.00 0.02 1.22 0.00
+MiniGPT-v2 33.06 37.52 34.72 32.56 9.81 8.28 84.95 85.91 26.98 25.94 36.76 38.64 60.42 80.72
+Otter 11.49 / 12.12 / 1.28 / 72.39 / $\emptyset$ / 35.33 / 75.88 /
+MiniGPT-4 34.11 / 15.05 / 1.88 / 87.09 / 30.20 / 37.19 / 67.37 /
+GPT-4o 25.81 / 18.67 / 2.00 / 66.67 / 21.68 / 40.54 / 43.06 /
+Gemini-2.5-flash 2.31 / 28.28 / 6.21 / 93.12 / 13.16 / 28.81 / 8.72 /
+Qwen-VL-2.5 19.13 / 15.34 / 2.70 / 75.55 / 10.63 / 33.85 / 36.70 /
+GLM-4.1-thinking 35.73 / 19.44 / 2.31 / 70.64 / 4.40 / 39.61 / 38.53 /
+Table 2: Benchmarking vision-language models on LEGO-VLM dataset.
+The upper part represents open-resourced vision-language models, both without (/wo) and with (/w) PEFT using LoRA.
+The lower part presents results obtained via API calls to the latest vision-language models.
 The bold font indicates the highest score in each column.
 Symbols $\uparrow$ and $\downarrow$ denote that higher and lower values are better, respectively.
-Symbol “–” indicates the model is not applicable for fine-tuning.
+Symbol “-” indicates the model is not applicable for fine-tuning.
 Symbol $\emptyset$” denotes a meaningless zero as the model fails to generate output as instructed.
@@
 We will explore this in future work, as this study primarily focuses on proposing the tasks rather than complex modeling.
-5.2 Data Quality Assessment
-To ensure the quality of the generated data for tasks T2 and T3, we sampled 100 data points and conducted quality assessments for each task.
+Data Quality Assessment
+To ensure the quality of the generated data for tasks T2 and T3, we sampled 100 data samples and conducted quality assessments for each task.
 We added bounding boxes based on the coordinates and asked three annotators to evaluate the generated data quality, focusing on coordinates, entities, and negative samples, based on the following criteria: (1) Entity disambiguitymeasures how clearly a target entity is relevant to the scene in T2, with scores of 0, 1, and 2 indicating low, medium, and high disambiguation, respectively.
@@
 (4) State identifiabilitymeasures whether the generated parts in an image are a recognizable variation of the original image in T3.
-Figure 4: Results of the quality assessment.
-As shown in Figure 4, the average entity disambiguation and boundary precision scores are 1.74 and 1.25 out of 2 with the Kappa values [^5] https://www.statsmodels.org/dev/generated/statsmodels.stats.inter_rater.fleiss_kappa.html of 0.60 and 0.61, respectively.
+Figure 4: Human assessment of data quality.
+As shown in Figure 4, the average entity disambiguation and boundary precision scores are 1.74 and 1.25 out of 2 with the Kappa values (Scikit-learn 2024) of 0.60 and 0.61, respectively.
+Query: [grounding] After I find 1 transparent mailbox front 2x2. Tell me what to do for the current situation. [Uncaptioned image]
+Reference of Assembly Instruction:
+Insert the two tabs at the bottom of the mailbox casing into the two small notches on the mailbox front.
+mPLUG-OWL2: Place the mailbox front 2x2 next to the block on the table   .
+BLIP2: Put the mailbox front 2x2 horizontally on the table, clasp to the back. Put the mailbox back 2x2 vertically on top, slope to the left.
+LlaVA: In the image, there is a small Lego-like object with a red liquid inside of it.
+The object is placed in a blue sky, and it appears to be a 3D model.
+To create a transparent mailbox, you need to find a transparent mailbox front that is 2x2 in size.
+Once you find it, you can use it as the base for your transparent mailbox.
+You can then place the small Lego-like object inside the mailbox to complete the scene.
+Qwen-VL: Put the transparent mailbox front 2x2 on the right side of the white arch.
+InstructBLIP: Place it horizontally on the table, clasp to the back.
+MiniGPT-v2: Here’s what you need to do with the pieces you have collected: 1.
+Place the transparent mailbox front 2x2 on top of the plate 1x2, aligned with the short side of the plate.
+Now you have completed the current situation.
+Let me know if you need help with anything else!
+Otter: In view of the reasoning ‘A mailbox is small and square with a see through window.’, the optimal answer should be: build structure.
+MiniGPT-4: Attach the transparent mailbox front to the base structure, aligning the tabs with the notches.
+GPT-4o: Ensure the transparent mailbox front 2x2 is securely attached by inserting the small shaft into the round brick 1x1 before proceeding with the next step.
+Table 3: An example for understanding VLM challenges.
 This suggests that annotators reached a moderate agreement on the clarity of the evaluated objects in relation to the corresponding scenes.
@@
 This ensures the overall quality of the generated states.
-5.3 Case Study
-Text Image
-Query [grounding] After I find 1 transparent mailbox front 2x2. Tell me what to do for the current situation. [Uncaptioned image]
-Reference Insert the two tabs at the bottom of the mailbox casing into the two small notches on the mailbox front.
-mPLUG-OWL2 Place the mailbox front 2x2 next to the block on the table.
-Fine-tuned Put it on the rightmost edge of the high part of the structure.
-BLIP2 Put the mailbox front 2x2 horizontally on the table, clasp to the back. Put the mailbox back 2x2 vertically on top, slope to the left.
-Fine-tuned Place the mailbox front 2x2 on the left side of the structure to complete the assembly.
-LlaVA In the image, there is a small Lego-like object with a red liquid inside of it.
-The object is placed in a blue sky, and it appears to be a 3D model.
-To create a transparent mailbox, you need to find a transparent mailbox front that is 2x2 in size.
-Once you find it, you can use it as the base for your transparent mailbox.
-You can then place the small Lego-like object inside the mailbox to complete the scene.
-Fine-tuned Put it on the mailbox.
-QWEN-VL Put the transparent mailbox front 2x2 on the right side of the white arch.
-Fine-tuned Put it on the table.
-InstructBlip Place it horizontally on the table, clasp to the back.
-Fine-tuned Put the mailbox front on the left side of the mailbox base.
-MiniGPT-v2 Sure!
-Here’s what you need to do with the pieces you have collected: 1.
-Place the transparent mailbox front 2x2 on top of the plate 1x2, aligned with the short side of the plate.
-Now you have completed the current situation.
-Let me know if you need help with anything else!
-Fine-tuned Put the mailbox front 2x2 horizontally in front of the previous piece.
-Otter In view of the reasoning ‘A mailbox is small and square with a see through window.’, the optimal answer should be: build structure.
-MiniGPT-4 Attach the transparent mailbox front to the base structure, aligning the tabs with the notches.
-GPT-4o Ensure the transparent mailbox front 2x2 is securely attached by inserting the small shaft into the round brick 1x1 before proceeding with the next step.
-Table 3: An example of scene understanding results.
-Words highlighted in orange represent manipulation verbs, red and blue indicate noun entities and their adjectives, and purple denotes positions.
-The word with a background color orange indicates the task-specific token and the words with a background color green is obtained from the previous step’s textual instruction manual.
+Case Study of Fine-grained VLM Challenges
 In this section, we present illustrative examples to highlight the capabilities and challenges in fine-grained vision-language understanding, particularly in achieving overlap with the reference and accurately recognizing entities and their properties, such as size, color, and etc.
 Table 3 present an example showcasing the intuitive results of the prevailing VLMs on scene understanding.
-We obtain the following observations: First, most models are able to correctly identify the core theme entity “mailbox front” (highlighted in green), with the exception of Otter and GPT-4o when performing scene understanding.
-However, generating a complete set of properties (e.g., “transparent” highlighted in blue and size “2x2” in pink) to describe these entities accurately and without ambiguity remains a significant challenge.
-Second, providing the correct assembly operations with precise positional instructions is highly challenging.
-Although LlaVA and MiniGPT-v2 generate informative instructions, they are not factually accurate in terms of the assembly process.
-6 Related Work
-6.1 Vision-Language Datasets in AR
-Dataset COIN HowTo100M TEACh MIMIC-IT HoloAssist EPIC-KITCHENS Assembly101 Cross-task RareAct LEGO-VLM (Ours)
-T1
-T2
-T3
-Size 11,827 23,611 3,215 2.8M 350 89,977 4,321 4,713 7,607 35,612
-Table 4: Comparison of datasets for assembly tasks towards AR training regarding the capabilities of tasks T1, T2, and T3. The symbols ✓and ✗indicate the presence or absence of each capability.
-Various datasets have been developed to enhance AR assembly tasks by supporting key capabilities T1, T2, and T3. A comparative summary is presented in Table 4.
-COIN provides a strong foundation for sequential task analysis with comprehensive annotations, aiding research in multimodal learning (Tang et al., 2019).
-HoloAssist captures egocentric human-AI interactions using mixed-reality headsets, offering valuable real-world insights (Wang et al., 2023b).
-HowTo100M, with its extensive collection of 136 million video clips and transcribed narrations, is highly effective for text-to-video retrieval but lacks fine-grained object annotations (Miech et al., 2019).
-TEACh focuses on interactive dialogues in domestic environments, improving dialogue modeling but not prioritizing detailed object detection (Padmakumar et al., 2021).
+Compared with the reference, existing VLM models face key challenges: (1) Vague and generic assembly instructions.mPLUG-OWL2 and BLIP2 produce broad, non-specific placement directives (e.g., “Place the mailbox front 2x2 next to the block on the table” or “Put the mailbox front 2x2 horizontally on the table, clasp to the back”), falling short of the precise, step-by-step guidance found in the reference.
+Otter’s response is even more superficial, merely stating “build structure” without actionable detail.
+(2) Struggles with fine-grained entity recognition and attribute grounding.While models like Qwen-VL and MiniGPT-v2 correctly mention the “transparent mailbox front 2x2,” others, such as InstructBLIP, overlook critical attributes (e.g., referring only to “it” or “the mailbox front”).
+LlaVA, despite its verbosity, hallucinates details (e.g., “red liquid inside” and “blue sky”) absent from the scene, highlighting confusion in entity identification and property association.
+(3) Tendency to hallucinate or misinterpret assembly steps.MiniGPT-v2 and GPT-4o introduce extraneous or incorrect actions, such as “aligning with the short side of the plate” or “inserting the small shaft into the round brick 1x1,” which are not part of the actual assembly process.
+This reveals a shallow grasp of both scene context and the logical flow of assembly operations.
+Overall, while some models can partially identify the correct entities and occasionally mention relevant properties, none consistently generate instructions that are both factually accurate and contextually appropriate for fine-grained assembly tasks.
+Related Work
+Vision-Language Datasets for Assembly Tasks
+Various datasets have been developed to enhance assembly tasks by supporting key capabilities T1, T2, and T3. A comparative summary is presented in Table 4.
+      Dataset       T1       T2       T3       Size
+      COIN       ✓       ✗       ✓       11,827
+      HowTo100M       ✓       ✗       ✓       23,611
+      TEACh       ✓       ✗       ✓       3,215
+      MIMIC-IT       ✓       ✗       ✗       2.8M
+      HoloAssist       ✓       ✗       ✓       350
+      EPIC-KITCHENS       ✗       ✓       ✓       89,977
+      Assembly101       ✓       ✗       ✓       4,321
+      Cross-task       ✗       ✓       ✓       4,713
+      RareAct       ✗       ✓       ✓       7,607
+      LEGO-VLM (Ours)       ✓       ✓       ✓       35,612
+Table 4: Compariable datasets for assembly tasks towards AR training regarding the cabablities of T1, T2, T3. The symbols ✓and ✗indicate the presence or absence of each capability.
+COIN provides a strong foundation for sequential task analysis with comprehensive annotations, aiding research in multimodal learning (Tang et al.
+2019).
+HoloAssist captures egocentric human-AI interactions using mixed-reality headsets, offering valuable real-world insights (Wang et al.
+2023b).
+HowTo100M, with its extensive collection of 136 million video clips and transcribed narrations, is highly effective for text-to-video retrieval but lacks fine-grained object annotations (Miech et al.
+2019).
+TEACh focuses on interactive dialogues in domestic environments, improving dialogue modeling but not prioritizing detailed object detection (Padmakumar et al.
+2021).
 While some datasets excel in action recognition, they often lack the procedural depth necessary for effective training.
-Assembly101, with over 4,000 toy assembly videos, does not include step-by-step instructional details (Sener et al., 2022).
-CrossTask facilitates weakly supervised learning by leveraging narrations and step lists but lacks temporal annotations to clarify action sequences (Zhukov et al., 2019).
-EPIC-KITCHENS provides extensive annotations on unscripted kitchen activities, yet it does not offer structured procedural guidance (Damen et al., 2018).
-RareAct presents unique interactions, challenging models to interpret complex actions without explicit instructions (Miech et al., 2020).
+Assembly101, with over 4,000 toy assembly videos, does not include step-by-step instructional details (Sener et al.
+2022).
+CrossTask facilitates weakly supervised learning by leveraging narrations and step lists but lacks temporal annotations to clarify action sequences (Zhukov et al.
+2019).
+EPIC-KITCHENS provides extensive annotations on unscripted kitchen activities, yet it does not offer structured procedural guidance (Damen et al.
+2018).
+RareAct presents unique interactions, challenging models to interpret complex actions without explicit instructions (Miech et al.
+2020).
 Stanescu et al.
 (2023a) introduce a state-aware prior that significantly improves object detection in assembly tasks like furniture and Lego construction.
-However, it lacks textual descriptions of object states and does not address fine-grained detection (Stanescu et al., 2023b).
-Meanwhile, MIMIC-IT contributes 2.8 million multimodal instruction-response pairs, enriching conversational modeling but lacking scenario-specific descriptions that would enhance task guidance (Li et al., 2023a).
+However, it lacks textual descriptions of object states and does not address fine-grained detection (Stanescu et al.
+2023b).
+Meanwhile, MIMIC-IT contributes 2.8 million multimodal instruction-response pairs, enriching conversational modeling but lacking scenario-specific descriptions that would enhance task guidance (Li et al.
+2023a).
 To bridge these gaps, we conduct a comparative analysis evaluating dataset coverage of essential components—objects, states, and scenes—critical for developing AR-guided assistants.
@@
 This dataset is designed to optimize the training and evaluation of advanced VLMs, improving their ability to handle fine-grained assembly tasks.
-6.2 Towards Fine-grained VLMs
-Recent advancements in vision-language models (VLMs) have broadened their applications, enhancing accessibility, search indexing, and interactive content retrieval (Fan et al., 2024; Zhang et al., 2024; Du et al., 2022; Gan et al., 2022).
-These models generate captions, describe images (Zhou et al., 2020; Hu et al., 2022), facilitate visual QA (Bazi et al., 2023).
-For instance, they can accurately identify objects in images based on phrases like “the red car” (Subramanian et al., 2022).
-Furthermore, these models facilitate conversational interactions about visual elements (Chen et al., 2022), which makes virtual assistants more adept at handling inquiries related to images and videos, which makes virtual assistants more adept at handling inquiries related to images and videos.
-Additionally, VLMs can support multilingual content (Gwinnup & Duh, 2023), benefiting assistive technologies and education (Chi et al., 2020; Wang et al., 2024a).
+Towards Fine-grained Vision-Lanuage Models
+Recent advancements in vision-language models (VLMs) have broadened their applications, enhancing accessibility, search indexing, and interactive content retrieval (Fan et al.
+2024; Zhang et al.
+2024; Du et al.
+2022; Gan et al.
+2022).
+These models generate captions, describe images (Zhou et al.
+2020; Hu et al.
+2022), facilitate visual QA (Bazi et al.
+2023).
+For instance, they can accurately identify objects in images based on phrases like “the red car” (Subramanian et al.
+2022).
+Furthermore, these models facilitate conversational interactions about visual elements (Chen et al.
+2022), which makes virtual assistants more adept at handling inquiries related to images and videos, which makes virtual assistants more adept at handling inquiries related to images and videos.
+Additionally, VLMs can support multilingual content (Gwinnup and Duh 2023), benefiting assistive technologies and education (Chi et al.
+2020; Wang et al.
+2024a).
 Despite progress, VLMs often lack fine-grained modeling for scene understanding, object recognition, and error detection.
-While CLIP demonstrates strong generalization (Radford et al., 2021), models like EfficientVLM (Wang et al., 2023a), MiniGPT-v2 (Chen et al., 2023), and Qwen-VL (Bai et al., 2023) struggle with intricate tasks.
-OSCAR (Li et al., 2020) and VisionLLM (Wang et al., 2024b) improve image-text alignment but face challenges in context-specific adaptation.
-The Otter model advances sequential task management yet falls short in detailed analysis (Li et al., 2023b).
-Efforts in procedural video representation (Zhong et al., 2023) and instructional task graphs (Ashutosh et al., 2024) highlight potential VLM integration but reveal gaps in error correction.
+While CLIP demonstrates strong generalization (Radford et al.
+2021), models like EfficientVLM (Wang et al.
+2023a), MiniGPT-v2 (Chen et al.
+2023), and Qwen-VL (Bai et al.
+2023) struggle with intricate tasks.
+OSCAR (Li et al.
+2020) and VisionLLM (Wang et al.
+2024b) improve image-text alignment but face challenges in context-specific adaptation.
+The Otter model advances sequential task management yet falls short in detailed analysis (Li et al.
+2023b).
+Efforts in procedural video representation (Zhong et al.
+2023) and instructional task graphs (Ashutosh et al.
+2024) highlight potential VLM integration but reveal gaps in error correction.
 To address these challenges, we propose a vision-language architecture optimized for fine-grained tasks, advancing training and benchmarking of VLMs.
-7 Conclusion
+Discussion
+Implications and limitations.
+This work addresses a critical gap in the development of intelligent, instruction-following multimodal assistants by focusing on fine-grained procedural understanding—an ability crucial for equitable access to education and skill development in a technologically evolving society.
+While our benchmark is partially synthetic and does not yet support 3D scene understanding, it represents an important step toward scalable, accessible AI systems that can interpret complex visual-textual instructions.
+The implications span beyond research: such systems could democratize hands-on learning by enabling visually impaired users to access assembly tasks through multimodal feedback, and could revolutionize technical education and vocational training by offering consistent, language-grounded support.
+Our findings further highlight the current limitations of VLMs in precise reasoning and verification, pointing toward future directions in responsible AI development that benefits society at large—including learners, workers, and underserved communities.
+Ethical Considerations.
+We recognize the ethical implications of developing VLMs for user interactions, so addressing these concerns is essential.
+To ensure ethical standards, we use open-source VLMs as benchmarks and prioritize user-centric design such as inclusiveness for diverse users.
+Conclusion
 Vision-language models (VLMs) have significantly advanced in recent years, enabling AI systems to interpret and generate textual descriptions of visual content.
 However, despite their success in general vision-language tasks, these models often struggle with fine-grained understanding, particularly in structured instructional scenarios.
-This research explores fine-grained vision-language modeling for manual-guide AR training, focusing on scene understanding, object detection, and state detection—areas where VLMs often struggle.
+This research explores fine-grained vision-language modeling for manual-guided LEGO assembly tasks, focusing on scene understanding, object detection, and state detection—areas where VLMs often struggle.
 We developed a specialized dataset from LEGO instruction manuals, designing fine-grained tasks to evaluate VLM performance in tracking assembly sequences and interpreting instructions.
@@
 Beyond technical advancements, our research aims to empower blind and visually impaired individuals by enabling AI-driven learning tools, promoting greater accessibility and independence.
-Appendix A Analysis
-A.1 Impact of Query Construction
-We explore the impact of query construction on the QWEN-VL model in the early phase to identify the optimal configuration for achieving the best inference results.
-Setting T1 T2 T3
-Instruction Example T_ACC $\uparrow$ ROUGE $\uparrow$ BLEU $\uparrow$ E_ACC $\uparrow$ IOU $\uparrow$ F1-State $\uparrow$ FPR $\downarrow$
-11.90 29.57 5.19 32.41 25.60 39.77 97.99
-6.84 24.63 4.58 5.56 22.01 34.00 97.50
-10.43 23.00 4.24 26.37 15.55 34.50 97.20
-5.27 11.02 1.30 3.70 0.10 25.50 98.40
-Table 5: A study of query construction results across various tasks and metrics.
-The symbol $\uparrow$ indicates higher is better, $\downarrow$ indicates lower is better, and values in bold represent the highest results performance.
-To this end, we conducted experiments on various query settings to determine which components most significantly influence performance.
-The model was chosen for this study due to its strong results under our base setting, which included both instruction directives and in-context examples.
-As shown in Table 5, our findings highlight the critical role of query design across different tasks.
-The results indicate that providing both instructions and examples yields the best overall performance, demonstrating their complementary roles in guiding inference.
-In the scene understanding task (T1), the presence of instructions plays a critical role in improving accuracy.
-When instructions are included, theme accuracy (T_ACC) increases significantly, suggesting that clear guidance helps the model better understand scene context and thematic elements.
-Furthermore, both ROUGE and BLEU scores, which evaluate the structure and coherence of generated outputs, improve when at least one of the two components (instructions or examples) is present.
-The lowest performance in this task occurs when neither component is provided, confirming that structured input is essential for meaningful inference.
-In the object detection task (T2), the results suggest that examples contribute more to entity recognition and output formatting than instructions alone.
-The entity accuracy (E_ACC) drops sharply from 32.41% to 5.56% when examples are removed, even when instructions remain.
-Similarly, IoU, which measures how well the output adheres to the expected format, significantly decreases in the absence of examples.
-This highlights that examples provide concrete reference points, enabling the model to generate well-structured and accurate entity labels.
-In the state detection task (T3), despite improvements in other metrics with enhanced query settings, the false positive rate (FPR) remains consistently high (above 97%) across all conditions.
-This suggests that the model has a strong affirmative bias, likely defaulting to generating positive responses regardless of query refinements.
-While instructions and examples improve structured outputs, they do not sufficiently mitigate the issue of excessive false positives.
-These findings emphasize the importance of query design in optimizing model performance.
-Combining instructions with instruction directives and in-context examples is crucial for achieving high accuracy, structured outputs, and reliable entity recognition.
-However, the persistent issue of high FPR suggests the need for further refinements, such as introducing negative examples, confidence calibration, or additional fine-tuning strategies to reduce overconfidence in affirmative responses.
-Future work should explore these approaches to enhance the model’s robustness and reliability in diverse inference tasks.
-A.2 Case Study of Object Detection
-Text Image
-Query [object] Please identify and highlight the positions of 1 light purple brick 2x6.
-Providing the positions in the format: <p>object</p> <Xleft><Ytop><Xright><Ybottom>, with X and Y coordinates normalized to [0,100].
-<Xleft> and <Ytop> for the top-left corner.
-<Xright> and <Ybottom> for the bottom-right corner.
-[Uncaptioned image]
-Reference <p>1 light purple brick</p> <58><15><92><35>
-mPLUG-OWL2 <p>1 light purple brick</p> <125><48><45><97>
-Fine-tuned <p>1 light purple brick</p> <60><24><69><35>
-BLIP2 <p>1 light purple brick</p> <57><15><82><35><p>and<p>1 dark purple brick 2x6</p> <82><25><93><36>
-Fine-tuned <p>1 light purple brick</p> <58><15><92><35>
-LLaVA The light purple brick is located at the center of the image, surrounded by yellow and black bricks.
-Fine-tuned <p>1 light purple brick</p> <62><15><87><35>
-QWEN-VL <20><61><30><78>
-Fine-tuned <p>1 light purple brick</p> <57><15><82><35>
-InstructiBLIP <p>1 light purple brick 2x6</p> <57><15><82><35>
-Fine-tuned <p>1 light purple brick</p> <58><15><92><35>
-MiniGPT-v2 <p>1 light purple brick</p> <62><20><88><40>
-Fine-tuned <p>1 light purple brick</p> <57><17><85><37>
-Otter The 1 light purple brick 2x6 is on the left side of the image.
-MiniGPT-4 Sure! Here are the positions of the light purple brick in the image: Xleft = 0, Ytop = 25, Xright = 30, Ybottom = 50.
-GPT-4o from PIL import Image # Load the image image_path $\dots$
-Table 6: An example of object detection results. The object is highlighted in light purple, and its position is highlighted in blue.
-Table 6 present an example of object detection task.
-We obtain the following observations: First, only InstructiBLIP and Otter generate the full description of the object “1 light purple brick 2x6” without disambiguation.
-Second, none of the models accurately predict the precise positions as outlined in the reference.
-Appendix B Enlightening Discussion
-B.1 Limitations
-The quality of object and state detection in our dataset heavily depends on the performance of current VLMs.
-While we have assessed the data quality, it is important to note that this is a synthetic dataset with only limited human intervention for corrections.
-In the object detection task, the model generates coordinates used to highlight detected objects with bounding boxes, though these coordinates are not meant for direct user interaction.
-In contrast, the outputs of the state detection and scene understanding tasks are text-based and designed to be user-friendly.
-However, we have not yet accounted for 3D scenes or objects.
-We evaluated a few existing VLMs as benchmarks, but their performance remains suboptimal.
-Future work will focus on improving fine-grained VLMs to enhance their effectiveness in these tasks.
-B.2 Broader Impact
-Our work has significant implications for accessibility and inclusive learning, particularly in assisting blind and visually impaired users (Gadiraju et al., 2020; Hao et al., 2024; Karamolegkou et al., 2024; Zhao et al., 2024).
-By integrating VLMs into conversational training assistants, we aim to create equitable learning opportunities in AR-based training environments.
-This could enable individuals with visual impairments to engage in hands-on assembly tasks through AI-powered guidance, reducing barriers to education and skill development.
-Beyond accessibility, our research contributes to industrial applications, particularly in manufacturing, robotics, and assembly training.
-Fine-grained vision-language reasoning can enhance AR-based instruction systems, reducing errors, improving efficiency, and enabling more intuitive human-AI collaboration in high-stakes environments like car manufacturing and industrial product assembly.
-By bridging the gap between VLMs and real-world AR applications, our work also fosters more intelligent and interactive training experiences, benefiting industries that rely on precise, step-by-step assembly guidance.
-B.3 Future Work
-A key area for future research is improving VLMs for fine-grained spatial reasoning.
-Current models struggle with detailed object and state recognition, limiting their ability to provide accurate real-time feedback in AR training.
-Advancements in pretraining strategies, multimodal fusion techniques, and architectural design could enhance spatial understanding in AR-based assembly tasks.
-Expanding the dataset is another critical direction.
-While our dataset focuses on LEGO-based assembly tasks as a simplified simulation, future work should explore real-world industrial assembly scenarios, such as furniture construction, electronic device assembly, and mechanical repairs.
-This would improve model generalization and practical applicability in professional training settings.
-Additionally, integrating real-time multimodal interaction is essential for developing dynamic AR assistants.
-Future research should explore VLMs that can process live user input, track assembly progress, and provide adaptive feedback based on visual and textual cues.
-This would move beyond static dataset-based evaluation toward fully interactive AR guidance systems.
-Finally, trust and explainability in human-AI collaboration remain crucial.
-Understanding how users interact with AI-powered AR assistants can help improve trust, usability, and transparency.
-Future work should investigate methods for making VLMs more interpretable and reliable, ensuring that AI-generated instructions align with human expectations and real-world assembly requirements.
-Appendix C Implementation Details
-C.1 Implementational setup
-During data synthesis, we utilized MiniGPT4-v2777https://github.com/Vision-CAIR/MiniGPT-4 as the VLM, employing straightforward rules and human correction to create vision-language data.
-To evaluate the performance of benchmark models, we randomly sampled 25% of the dataset to serve as the test set.
-All inference and fine-tuning experiments were conducted on NVIDIA A100 SXM4 40GB GPUs.
-We set the learning rate to 1e-5 and the batch size to 16 for our experiments.
-The AdamW optimizer (Loshchilov & Hutter, 2017) was employed to facilitate efficient training, which spanned 3 epochs.
-C.2 Query Templates
-In this paper, we investigate three fine-grained vision-language tasks and have devised specific query templates for each, as demonstrated in Table 7.
-T1: Scene Understanding
-Token: [grounding]
-Selection Rule: <step_symbol> with <step>. Instruction from this step.
-Format-Directive: Tell me what to do for the current situation.
-T2: Object Detection
-Token: [object]
-Selection Rule: <step_symbol> with <eop>, Word starts with “find” or “collect”. Instruction from this step.
-Format-Directive: Providing the positions in the format: <p>object</p> {<|Xleft><Ytop><Xright><Ybottom>}, with X and Y coordinates normalized to [0,100].
-<Xleft> and <Ytop> for the top-left corner.
-<Xright> and <Ybottom> for the bottom-right corner.
-T3: State Detection
-Token: [state]
-Selection Rule: <step_symbol> with <step>. Previous <step_symbol> is <eop>. Instruction from this step and the previous step.
-Format-Directive: Just tell me Yes or No.
-Table 7: Query Templates for Fine-Grained VLM Tasks.
-C.3 Dataset Details
-C.3.1 Data Schema in JSON Format
-This section presents the JSON data schema for integrating multimodal data (visual and textual) and fine-tuning VLM.
-The schema defines key components that structure and organize data efficiently, ensuring seamless preprocessing before final compilation.
-  "instruction_id": 95,
-  "text": [ "Find 2 transparent red flat tiles 1x2." ],
-  "VLM": {
-    "img_path": "...",
-    "parts_img_path": "None",
-    "id_sign": "Car",
-    "step_num": 28,
-    "step_class": "eop",
-    "other_sign": "AG",
-    "background_rect": "None",
-    "color_list": [],
-    "bound_color": [],
-    "task_label": "object",
-    "query": "Find 2 transparent red flat tiles 1x2.",
-    "MiniGPTv2_output": "<p>2 transparent red flat tiles</p> {<28><17><73><50>}"
-  "img": "...",
-  "entities": [
-    { "this_line": "Find 2 transparent red flat tiles 1x2." },
-    { "verb": "Find" },
-    { "det": "2" },
-    { "colorinfo": "transparent red" },
-    { "theme": "flat tiles 1x2" },
-    { "suffix": "." } ]
-Table 8: An example of a JSON schema for integrated multimodal data.
-The key fields related to VLM in the JSON schema include:
-  • img_path: Stores the local path of images downloaded via web scraping, pointing to the original files fetched from the web.
-  • id_sign: Extracted from the image filename, this field identifies the sub-assembly session (e.g., “CAR” in this example).
-  • step_num: Represents the sequence number of the assembly step (e.g., step 28), tracking the assembly’s progression.
-  • step_class: A classification tag (e.g., “eop” for “end of part”) indicating an object-finding step in the assembly process.
-  • MiniGPTv2_output: Stores processed output from MiniGPTv2, containing annotations or modifications essential for subsequent dataset creation.
-  • parts_img_path, bound_color, color_list: Reserved for assembly steps classified as “step” and crucial for state detection dataset generation:
-    • color_list: Records all possible boundary colors used throughout the assembly session.
-    • bound_color: Consolidates the most probable boundary colors through a voting mechanism, ensuring consistency across the dataset.
-To support fine-tuning, our default JSON format follows the Llava model’s fine-tuning data structure, ensuring an optimized and adaptable framework for general use cases.
-An example of the JSON schema for VLM fine-tuning is illustrated in Table 9.
-  "id": "lego-60263-ocean-mini-submarine-readscr-12",
-  "image": "sm01-island-0_step_Step%20%23%23E.png",
-  "conversations": [
-      "from": "human",
-      "value": "[grounding] After I find 1 reddish brown chest bottom 2x4.
-      Tell me what to do for the current situation."
-      "from": "gpt",
-      "value": "Put it horizontally on the table."
-Table 9: An example of a JSON schema for VLM fine-tuning.
-The full dataset, comprising JSON and PNG files, occupies approximately 11.8 GB. For specialized requirements, such as those for the Otter model, we convert all text and image data into Parquet files.
-C.3.2 An Example of LEGO Instruction Manual
-Figure 5 shows an example of a part of the instruction manual we crawled from the official LEGO website.
-Figure 5: An example of LEGO instruction manual.
-It consists of a summary section at the beginning, followed by three sequential instruction steps.
-Each step includes textual instructions paired with corresponding images to guide the assembly process.
-The LEGO audio instruction manual is a structured guide designed to assist blind and visually impaired users in assembling LEGO sets accurately.
-The key sections of a LEGO instruction manual include: (1) a summary providing general information, such as topics and candidate parts for assembly, followed by (2) a sequence of multimodal step instructions.
-Each step consists of textual instructions paired with an illustrative image, with key functional phrases, such as theme entities, highlighted for clarity provided by LEGO.
-With an average of 215.3 steps per manual, we segmented each manual into multiple assembly sessions, each consisting of a paired sequence: an object-seeking step followed by an object-assembly step.
-C.3.3 Illustration of Data Creation.
-Figure 6 illustrates the process of creating the dataset for task T2.
-It begins by iterating through object-seeking steps, where users locate specific objects.
-Next, images are processed to generate the object detection dataset, consisting of query-image-text triplets.
-Queries are formed using a template with the task token "[object]" and the current instruction.
-Responses include positional text indicating object coordinates, initially inferred using MiniGPT-v2.
-A composite image is then created by merging the target image with three randomly sampled images, and the coordinates are adjusted accordingly.
-Figure 6: An illustration of object detection data creation.
-Figure 7 illustrates the process of generating the dataset for task T3.
-The procedure begins by iterating through object-assembly steps, pairing each with its preceding step where users identified parts to assemble.
-Next, part segmentation is performed by detecting boundary colors using K-means clustering and Hue filtering, followed by manual correction.
-Boundaries are highlighted in distinct colors to mark assembly objects.
-Bounding boxes are then added to segmented objects, and natural perturbations are introduced by shifting assembled parts within the background box.
-Each correct assembly state serves as a positive sample, while three perturbed states act as negative samples.
-Finally, the dataset is constructed, where queries integrate task tokens, current and previous instructions, and responses indicate correct or incorrect assembly states.
-Figure 7: An illustration of synthesized negative samples for state detection.
-C.3.4 Examples in Data Quality Assessment
-Figure 8 and 9 illustrate examples of the quality assessment for generated data in object detection (T2) and state detection (T3), respectively.
-We randomly selected 100 data examples per task to ensure a fair assessment.
-Three annotators were provided with comprehensive guidelines on how to evaluate each dataset based on multiple criteria, accompanied by relevant examples.
-The first five data points mentioned in the guidelines were used as a sanity check to ensure that the annotators correctly understood and applied the evaluation criteria.
-If annotators fail the sanity check, they undergo retraining until they pass.
-We evaluate annotations using the average scores of four criteria (i.e., entity disambiguity, boundary precision, state relevance, and state identifiability) and calculate Kappa values to assess annotation agreement.
-Figure 8: An example of evaluating data quality of task T2.
-Figure 9: An example of evaluating data quality of task T2.